factor 2x^2+12=x^2+8x
Answer by jim_thompson5910(35256)

2x^2+12=x^2+8x

2x^2+12-x^2-8x = 0

x^2-8x+12 = 0

(x-6)(x-2) = 0

x-6 = 0 or x-2 = 0

x = 6 or x = 2

So the solutions are x = 6 or x = 2